# Amortized Bayesian Meta-Learning for Low-Rank Adaptation of Large Language Models

### v2

@@ A.1 Additional Experiment Results
A.2 Experiment Details
Here we detail the experimental setup used in our experiments. We use PyTorch with Torchtune to fine-tune Llama3-8B-Chat. Each experiment uses a single A100 GPU with 40GB memory. Each experiment uses batch-size of 2 and 5 meta-learning adaptation steps. For meta-learning methods, inner loop size of 5 is used.
We detail hyperparameters in Table 5. All methods use the same LoRA setup, which is detailed in 6.
A.3 Datasets
We use the setup of ye-etal-2021-crossfit for our cls-45 and cls-23 datasets (the Winogrande one uses the same training set as cls-45). We utilized the codebase provided by min-etal-2022-metaicl to setup the datasets. Additionally, we focus on problems that can be converted into the multiple choice format. This allows us to evaluate the calibration error of models. Filtering for questions with at most four choices, we get the following training, validation, and test splits of these datasets.
@@ A.3 Datasets
cls-23 testing: same as cls-45 testing.
@@
As for the validation versus testing splits of other datasets, we follow the splits provided by min-etal-2022-metaicl.

### v3

@@ A.2 Additional Experiment Results
A.3 Experiment Details
Here we detail the experimental setup used in our experiments. We use PyTorch with Torchtune to fine-tune Llama3-8B-Chat. Each experiment uses a single A100 GPU with 40GB memory. All methods use the AdamW optimizer (Loshchilov and Hutter, 2017), batch-size of 2, inner-loops with 5 gradient steps, LoRA adapters with rank =8=8 following the standard practice, and learning rate is tuned in $10^{-6},10^{-4}]$}. For ABMLL, $\beta=10^{-8}$}, $c=e^{-20}$}. For the gamma prior, $a_{0}=1,b_{0}=0.01$} following Ravi and Beatson (2019).
We detail hyperparameters in Table 7. All methods use the same LoRA setup, which is detailed in Table 8.
A.4 Datasets
We use the setup of Ye et al. (2021) for our cls-45 and cls-23 datasets. We utilized the codebase provided by Min et al. (2022) to setup the datasets. Additionally, we focus on problems that can be converted into the multiple choice format. This allows us to evaluate the calibration error of models. Filtering for questions with at most four choices, we get the following training, validation, and test splits of these datasets.
@@ A.4 Datasets
cls-23 validation testing: same as cls-45 validation testing.